OpenStudy (anonymous):

the general term of a sequence is given. Determine whether the given sequence is arithmetic, geometric, or neither. If the sequence is arithmetic, find the common difference; if it is geometric, find the common ratio.

OpenStudy (anonymous):

\[an = 4n -2\]

OpenStudy (anonymous):

I think it is geometric

OpenStudy (anonymous):

lets take each value of n... a0=4.0-2=-2 a1=4.1-2=2 a2=4.2-2=6 a3=4.3-2=10 and so on.. hence the sequence is in AP and the common difference is 4 :)

OpenStudy (anonymous):

oh so you just place a value in n and solve when one is not given

OpenStudy (anonymous):

yes...... thats the simple technique for such questions :)))))) @moser90

OpenStudy (anonymous):

thank you